I have to find where the parabola crosses the x-axis for y=x(3-x). I know the answer is x=0 and x=3 but I don't understand why it is not x= -3. Could somebody please explain this to me?
Answer by Alan3354(69443)

y = x(3-x) = 0
x*(3-x) = 0
--> either x = 0, or
3-x = 0
x = 3